OpenStudy (anonymous):

How would u rewrite y=a(x-h)^2

OpenStudy (anonymous):

you mean expand it?

OpenStudy (anonymous):

\[y=ax ^{2}-2axh+ah ^{2}\]

OpenStudy (anonymous):

No like how would you use that form in this equation y=x^2-6x

OpenStudy (anonymous):

you mean in a form of y=x²+x ?

OpenStudy (anonymous):

No I mean in this form y=a(x-h)^2 with this problem y= x^2 - 6x how would you find the vertex of that problem.

OpenStudy (anonymous):

You are not giving me enough information sorry, can you please type out the whole question?

OpenStudy (anonymous):

How would you find the vertex of this problem y= x^2 -6x and rewrite in y=a(x-h)^2 form

OpenStudy (anonymous):

Y=a(x-h)^2 + k my bad